A gas phase reaction at equilibrium has the equilibrium expression: (PCS2 ) x (PH2)4 / (PCH4) x (PH2S)2 What is the reaction equation?

A. CH4 + H2S ⇌ CS2 + H2

B. CS2 + 4 H2 ⇌ CH4 +2 H2S

C. CH4 + 2 H2S ⇌ CS2 + 4 H2

D. 2 CH4 + H2S ⇌4 CS2 + H2

E. CS2 + CH4 ⇌ H2 + H2S
Answer #1

A. CH4 + H2S ⇌ CS2 + H2

   Kp = (PCS2)(PH2)/(PCH4)(PH2S)

B. CS2 + 4 H2 ⇌ CH4 +2 H2S

   KP   = (PCH4)(PH2S)^2/(PCS2)(PH2)^4

C. CH4 + 2 H2S ⇌ CS2 + 4 H2

    KP = (PCS2)(PH2)^4/ (PCH4)(PH2S)^2

D. 2 CH4 + H2S ⇌4 CS2 + H2

    Kp   = (PCS2)^4(PH2)/ (PCH4)^2(PH2S)

E. CS2 + CH4 ⇌ H2 + H2S

   Kp   = (PH2)(PH2S)/(PCS2)(PCH4)

C. CH4 + 2 H2S ⇌ CS2 + 4 H2 >>>>>answer

    KP = (PCS2)(PH2)^4/ (PCH4)(PH2S)^2